OpenStudy (serenavdw):

simplify each expression (6x^2-8x-7x^3)-(6x^3-x-6)

OpenStudy (serenavdw):

the x^2 or x^3 means x to the second or third power

OpenStudy (anonymous):

I'm pretty sure it's -17x-11

OpenStudy (serenavdw):

how did you get it?

OpenStudy (anonymous):

you add the x's together so 6+6=12 and -7+-7+-7= -21, 12x-8x-21x=-17x then you do the other half 6+6+6= 18x-x this x right here has a imaginary 1 so it would be 18x-1x-6. opps the answer is actually -17x-17x-6 sorry

OpenStudy (anonymous):

becuase you dont add the 6 in because it has no x next to it

OpenStudy (anonymous):

18x-1x=17x so altogether your answer is -17x-17x-6

OpenStudy (anonymous):

do you get it? @serenavdw

OpenStudy (akashdeepdeb):

\[(6x^2-8x-7x^3)-(6x^3-x-6)\] Now let us take out the parenthesis. They make things confusing! Now to take them out, apply distributive laws. [-(a+b+c) = -a -b -c] \[6x^2-8x-7x^3-6x^3+x+6\] Now add all like terms and what I mean by that is add the co-efficients of the like degree polynomials. For example. [A generalized example] I have \[ax^n + bx^n + cx^{n-1} + dx^{n-1} + ex^n\] Let us add the like terms with the same \(~~x^{anything}\) Which'll give us:\[(a+b+e)x^n + (c+d)x^{n-1}\] Coming back to the question: Adding the co-efficients of the like terms:\[6x^2 + (-8 + 1)x + (-7-6)x^3 + 6\] \[6x^2 - 7x - 13x^3 + 6\] Following mathematical convention we write this down with decreasing order of the power on 'x'. \[-13x^3 + 6x^2 - 7x + 6\] Getting this? :)

OpenStudy (serenavdw):

i got what akashdeepDeb got not yours

OpenStudy (serenavdw):

thanks to both of you though :)